OpenStudy (anonymous):

HALP!!! It's not an open ended question. Find g(f(3)). f(x) = x^3, g(x) = 4x + 3

geerky42 (geerky42):

Can you find f(3)?

geerky42 (geerky42):

\[\Large f(x)=x^3\\~\\\Large f(3)=3^3=\cdots~?\]

OpenStudy (confluxepic):

First find f(3) by plugging in the value. Then find g(f(x)) by plugging in the value.

OpenStudy (confluxepic):

I am not sure after that.

OpenStudy (confluxepic):

Good luck.

geerky42 (geerky42):

You are correct

OpenStudy (confluxepic):

Thanks.

OpenStudy (anonymous):

I got 405. Is that right? The choice is on the test.

geerky42 (geerky42):

How? I got 111.

geerky42 (geerky42):

\(3^3=27\), right? So \(f(3) = 3^3 = 27\) So \(g(f(3)) = g(27)=4(27)+3=\cdots~?\)

OpenStudy (anonymous):

405 was correct. yay

geerky42 (geerky42):

Apparently you wrote f(x) and g(x) incorrectly.

OpenStudy (anonymous):

erm...ok

OpenStudy (anonymous):

thnx btw